# SASSL: Enhancing Self-Supervised Learning via Neural Style Transfer

### v2

@@ 5.3 Ablation Study
Downstream training settings. For the ablation study, we cover four scenarios: (i) default MoCo v2, (ii) SASSL $+$ MoCo v2 using only in-batch representation blending, i.e., without interpolation ($\beta=1$), (iii) SASSL $+$ MoCo v2 using in-batch representation blending and pixel interpolation, and (iv) SASSL $+$ MoCo v2 using all its attributes (representation blending, pixel interpolation and an external style dataset).
@@
Results. Table 3 shows our ablation study on MoCo v2. It highlights the importance of controlling the amount of stylization using both representation blending and image interpolation. Without image interpolation, using style transfer as augmentation degrades the downstream classification performance by more than $1.5\%$ top-1 accuracy.
@@
On the other hand, by balancing the amount of stylization via blending and interpolation, SASSL boosts performance by more than $1.5\%$. This is a significant improvement for the challenging ImageNet scenario. Finally, by incorporating an external dataset such as Painters by Numbers, we further improve downstream task performance by almost $2.1\%$ top-1 accuracy. This shows the importance of diverse style references and their effect on downstream tasks.
Please refer to Section A.7 for additional ablation studies on the effect of the style latent representation ${\bm{z}}_{s}$ and the number of stylized layers on the downstream accuracy.
@@ 6 Conclusion
We propose SASSL, a novel data augmentation approach based on Neural Style Transfer that exclusively transforms the style of training samples, diversifying data augmentation during pre-training while preserving semantic attributes. We empirically show our approach outperforms well-established methods such as MoCo v2 by more than $2\%$ top-1 classification accuracy on the challenging ImageNet dataset. Importantly, SASSL also improves the transfer capabilities of learned representations, enhancing linear probing and fine-tuning classification performance across image domains by more than $3\%$ and $1\%$ top-1 accuracy, respectively. Future work includes extending our technique to different SSL methods and encoder architectures (e.g. ViTs), which can readily be combined with our method.
7 Acknowledgements
@@ Appendix A Appendix
The appendix is organized as follows
  • Sec. A.1 shows the details of the target and style datasets used in our downstream evaluations.
  • Sec. A.2 provides insight on how to select an external style dataset by analyzing their style representations.
  • Sec. A.3 includes additional information on the pre-training settings used in our experiments.
  • Sec. A.4 includes detailed information on the downstream task settings used in our experiments.
  • Sec. A.5 shows additional experiments on downstream performance using SimCLR and BYOL.
  • Sec. A.6 reports few-shot classification performance of MoCo v2 equipped with SASSL.
  • Sec. A.7 includes ablation studies to better understand the effect of each component of Neural Style Transfer (NST) in the downstream performance.
  • Sec. A.8 covers the computational requirements of our proposed method.
  • Sec. A.9 covers recent work on semantic-aware data augmentation and NST, and discusses their differences with our proposed method.
  • Sec. A.10 shows additional experiments on downstream performance using a more complex representation model (ResNet-50 x4) and an alternative representation architecture (ViT-B$/16$).
@@ A.1 Target and Style Datasets
We provide the details of the image datasets used in our experiments. Table 4 covers both target and style datasets, including their size and splits.
Table 4: Target and Style Datasets. Additional details on data split and number of samples of the image datasets used in our experiments.
Dataset Task Train Split Val. Split Test Split
ImageNet Pre-training, Target, Style $1,281,167$- $50,000$
ImageNet $1\%$ Target $12,811$- $50,000$
i_Naturalist 2021 Target, Style $2,686,843$- $500,000$
Diabetic Retinopathy Detection Target, Style $35,126$ $10,906$ $42,670$
Describable Textures Dataset Target, Style $1,880$ $1,880$ $1,880$
Painter by Numbers Style $79,433$- $23,817$
@@ A.2 Style Dataset Selection
Our transfer learning results in Section 5 demonstrate that SASSL achieves improved or comparable downstream performance across multiple datasets by incorporating NST as data augmentation. This raises an important question: how can we select an external style dataset to ensure downstream accuracy improvement? Here, we delve into the similarity between datasets in terms of their styles and establish its connection to the performance improvement gained by using them as external style references.
We focus on the linear-probing scenario as it freezes the representation model, forcing the classification head to rely on the representations learned during pre-training (rather than updating them as is the case with fine-tuning) to distinguish between the target categories.
As an example, in our linear probing experiments presented in Table 2, when Diabetic Retinopathy is used as the target dataset, the downstream accuracy achieved via SASSL + MoCo v2 is comparable to that of the default MoCo v2 algorithm, meaning there is no improvement in performance. We hypothesize that this is because the style representations of Diabetic Retinopathy are significantly different from those of the pre-training (ImageNet) and style datasets. Therefore, learning representations that are invariant to such a distinct set of styles does not contribute to distinguishing between target classes.
To support our hypothesis, we visualize the relationship between style representations using low-dimensional embeddings generated via t-SNE (Van der Maaten & Hinton, 2008) to capture the similarity between styles of different datasets. Style representations of each dataset, corresponding to vectors of length $100$, are obtained using an InceptionV3 feature extractor, as done by the FastSt algorithm. Next, we randomly select 1,800 style representations from each class and compute their two-dimensional embeddings using a perplexity of $30$, early exaggeration of $12$, and initializing the dimensionality reduction using PCA. We compute embeddings using 2,048 iterations.
Figure 4 depicts the low-dimensional embeddings obtained via t-SNE from all datasets used in our transfer learning experiments. The low-dimensional representation shows that the style representations from Diabetic Retinopathy are significantly distinct from those of the rest of datasets, including ImageNet. This aligns with our hypothesis, suggesting that SASSL improves transfer learning when the pre-training and style references are similar to those of the target dataset. From the perspective of t-SNE embeddings, this implies that pre-training and style datasets must have a good overlap with the target dataset for SASSL to improve downstream performance.
@@ A.3 Additional Experimental Details
All of our experiments were implemented in Tensorflow (Abadi et al., 2016). All our models were pre-trained on $64$ TPUs using a batch size of $4,096$. Both the MoCo v2 models pre-trained using the default augmentation and our proposed SASSL approach do not use a dictionary queue.
@@
When using SASSL, we applied style transfer after random cropping and before random horizontal flipping. Our default hyperparameters for SASSL were blending and interpolation factors drawn randomly from a uniform distribution between $0.1$ and $0.3$, and a probability of $0.8$ that style transfer will be applied.
@@ A.4 Downstream Training and Testing Settings
A.5 Downstream Task Performance on other SSL Methods
We extend the downstream performance experiments reported in Sec. 5.1 by evaluating SASSL’s impact on two other well-established SSL methods: SimCLR and BYOL. We pre-trained a ResNet-50 backbone using each method, both with and without SASSL, and then evaluated the resulting models on the ImageNet dataset using linear probing. In both cases, we employed the default pre-training and linear probing configurations provided by each method. SASSL was implemented with its recommended hyperparameters (blending and interpolation factor between $0.1$ and $0.3$, probability $0.8$, and Painter by Numbers as external style dataset). The pre-training and fine-tuning settings for SimCLR and BYOL adhered to their default configurations.
Table 5 shows the top-1 and top-5 downstream classification accuracy. Results shows that SASSL consistently improves the downstream classification accuracy of both SimCLR and BYOL, suggesting its effectiveness across different SSL techniques.
Table 5: SimCLR and BYOL Downstream classification performance on ImageNet. Classification accuracy of SimCLR and BYOL equipped with our proposed style transfer augmentation approach (SASSL). Top-1 and top-5 accuracy reported on a single random seed.

### v4

@@ 5.5 Ablation Studies
SASSL components. For the ablation study, we cover four cases: (i) MoCo v2 with default augmentation, (ii) SASSL + MoCo v2 using in-batch representation blending and no pixel interpolation ($\beta=1$), (iii) SASSL + MoCo v2 using in-batch representation blending and pixel interpolation, and (iv) SASSL + MoCo v2 using all its attributes (blending, interpolation and an external style dataset).
@@
Table 5.5 shows our ablation study using MoCo v2 as SSL technique. Results highlight the importance of controlling the amount of stylization using both representation blending and image interpolation. Without image interpolation, using Style Transfer as data augmentation degrades the downstream classification performance by more than $1.5\%$ top-1 accuracy.
@@
On the other hand, by balancing the amount of stylization via blending and interpolation, SASSL boosts performance by more than $1.5\%$. This is a significant improvement for the challenging ImageNet scenario. Finally, by incorporating an external style dataset such as PBN, we further improve downstream task performance by almost $2.1\%$ top-1 accuracy. This shows the importance of diverse style references and their effect on downstream tasks.
[…]
Number of stylized layers. We explore how the number of layers used to apply style transfer via CIN affects downstream performance. We analyze three cases: (i) stylizing using the first two residual blocks of the Stylization Network $\mathcal{T}$ (four layers from blocks 1 and 2), (ii) the first four residual blocks (eight layers from blocks 1 to 4), and (iii) all five residual blocks (ten layers).
For each case, we pretrain and linearly probe a ResNet-50 on ImageNet using SASSL + MoCo v2 with its recommended settings ($\alpha,\beta\in[0.1,0.3]$, $p=0.8$) and PBN as style dataset. To fully remove the effect of a style embedding ${\bm{z}}_{s}$, our comparison includes a model pretrained using the content image itself as style reference ($\hat{{\bm{z}}}={\bm{z}}_{c}$). We also compare our full SASSL + MoCo v2 model, stylizing all residual and upsampling blocks of $\mathcal{T}$ (thirteen layers).
Table 5.5 shows a progressive enhancement in accuracy with increasing stylization depth. Adding stylization to the first four layers showed negligible gains, mirroring the accuracy of the unaligned model. Stylizing the first eight and ten layers yielded modest improvements of $0.34\%$ and $0.52\%$, respectively, implying a growing influence of deeper layers on accuracy. Notably, pretraining with full stylization, encompassing both residual and upsampling layers, attains a $1.61\%$ accuracy boost, suggesting the importance of aligning deeper upsampling layers for downstream performance.
@@ 6 Conclusion
We propose SASSL, a novel data augmentation approach based on Neural Style Transfer that exclusively transforms the style of training samples, diversifying data augmentation during pretraining while preserving semantic attributes. We empirically show our approach outperforms well-established methods such as MoCo v2, SimCLR and BYOL by up to $2\%$ top-1 classification accuracy on ImageNet. SASSL also improves the transfer capabilities of learned representations, enhancing linear probing and fine-tuning performance across domains by up to $10\%$ and $6\%$ top-1 accuracy, respectively. Our technique can be extended to other SSL methods and models with minimum hyperparameter changes, as experimentally shown.
Broader Impact Statement
This work proposes a novel data augmentation approach leveraging Neural Style Transfer to enhance Self-supervised Learning, particularly for domains with limited data or expensive annotations. Our method utilizes semantic-aware image preprocessing to extract robust representations that generalize across diverse domains. This advancement tackles the critical challenge of using unlabeled data for Deep Learning, which has many potential positive impacts in both technical and societal fronts. SASSL’s style transfer component helps to reduce sensitivity to image texture, potentially improving model robustness to texture bias. However, since our method primarily modifies data augmentation, it may not fully address other potential biases arising from pre-training datasets or learning strategies.
Acknowledgements
@@ Appendix A Appendix
The Appendix is organized as follows:
  • Section A.1 shows the details of the target and style datasets used in our downstream evaluations.
  • Section A.2 provides insight on how to select an external style dataset by analyzing their style representations.
  • Section A.3 includes additional information on the pretraining settings used in our experiments.
  • Section A.4 includes detailed information on the downstream task settings used in our experiments.
  • Section A.5 reports full downstream classification accuracy (mean and standard deviation) on our SASSL + MoCo v2 experiments.
  • Section A.6 includes an additional ablation study to better understand the effect of Neural Style Transfer in the downstream performance.
  • Section A.7 covers the computational requirements of our proposed method.
@@ A.1 Target and Style Datasets
We provide the details of the image datasets used in our experiments. Table 2 covers both target and style datasets, including their size, splits and number of classes.
Table 2: Target and Style Datasets. Additional details on number of classes, data split and samples of the image datasets used in our experiments.
Dataset Task Classes Train Split Val. Split Test Split
ImageNet Deng et al. (2009) Pretraining, Target, Style $1,000$ $1,281,167$- $50,000$
iNaturalist ‘21 (iNaturalist, 2021) Target, Style $10,000$ $2,686,843$- $500,000$
Diabetic Retinopathy Detection (Kaggle & EyePacs, 2015) Target, Style $5$ $35,126$ $10,906$ $42,670$
Describable Textures Dataset (Cimpoi et al., 2014) Target, Style $47$ $1,880$ $1,880$ $1,880$
Painter by Numbers (Kan, 2016) Style $1,584$ $79,433$- $23,817$
ImageNet $1\%$ (Chen et al., 2020a) Target $1,000$ $12,811$- $50,000$
Food101 (Bossard et al., 2014) Target $101$ $75,750$- $25,250$
CIFAR10 (Krizhevsky, 2009) Target $10$ $50,000$- $10,000$
CIFAR100 (Krizhevsky, 2009) Target $100$ $50,000$- $10,000$
SUN397 (Xiao et al., 2010) Target $397$ $76,128$ $10,875$ $21,750$
Cars (Krause et al., 2013) Target $196$ $8,144$- $8,041$
Caltech-101 (Fei-Fei et al., 2004) Target $102$ $3,060$- $6,084$
Flowers (Nilsback & Zisserman, 2008) Target $021$,02010201,0201 , 020 $1,020$ $6,149$
@@ A.2 Style Dataset Selection
Our transfer learning results in Section 5 demonstrate that SASSL achieves improved or on-par downstream performance across multiple datasets by incorporating Neural Style Transfer (NST) as data augmentation. This raises an important question: how can we select an external style dataset to ensure downstream accuracy improvement? Here, we delve into the similarity between datasets in terms of their styles and establish its connection to the performance improvement gained by using them as external style references.
We focus on the linear-probing scenario as it freezes the representation model, forcing the classification head to rely on the representations learned during pretraining (rather than updating them as is the case with fine-tuning) to distinguish between the target categories.
As an example, in our linear probing experiments presented in Table 1, when Diabetic Retinopathy is used as the target dataset, the downstream accuracy achieved via SASSL + MoCo v2 is comparable to that of the default MoCo v2 algorithm, meaning there is no improvement in performance. We hypothesize that this is because the style representations of Diabetic Retinopathy are significantly different from those of the pretraining (ImageNet) and style datasets. Therefore, learning representations that are invariant to such a distinct set of styles does not contribute to distinguishing between target classes.
To support our hypothesis, we visualize the relationship between style representations using low-dimensional embeddings generated via t-SNE (Van der Maaten & Hinton, 2008) to capture the similarity between styles of different datasets. Style representations of each dataset, corresponding to vectors of length $100$, are obtained using an InceptionV3 feature extractor, as done by the Fast Style Transfer algorithm. Next, we randomly select 1,800 style representations from each dataset and compute their two-dimensional embeddings using a perplexity of $30$, early exaggeration of $12$, and initializing the dimensionality reduction using PCA. We compute embeddings using 2,048 iterations.
Figure 2 depicts the low-dimensional embeddings obtained via t-SNE from six diverse datasets used in our transfer learning experiments. The low-dimensional representation shows that the style embeddings from Diabetic Retinopathy are significantly distinct from those of the rest of datasets, including ImageNet. This aligns with our hypothesis, suggesting that SASSL improves transfer learning when the pretraining and style references are similar to those of the target dataset. From the perspective of t-SNE embeddings, this implies that pretraining and style datasets must have a good overlap with the target dataset for SASSL to improve downstream performance.
@@ A.3 Additional Experimental Details
All of our experiments were implemented in Tensorflow (Abadi et al., 2016). All our models were pretrained on $64$ TPUs using a batch size of $4,096$. Both the MoCo v2 models pretrained using the default augmentation and our proposed SASSL approach do not use a dictionary queue.
@@
When using SASSL, we applied NST after random cropping and before random horizontal flipping. Our default hyperparameters for SASSL were blending and interpolation factors drawn randomly from a uniform distribution between $0.1$ and $0.3$, and a probability of $0.8$ that NST will be applied.
@@
Table 3: SASSL $+$ MoCo v2 downstream classification accuracy on ImageNet. Linear probing accuracy of ResNet-50 pretrained via SASSL $+$ MoCo v2. Mean and standard deviation reported over five random trials.
[…]
MoCo v2 (Default) $72.55\pm 0.67$ $91.19\pm 0.34$
SASSL $+$ MoCo v2 (Ours) $\bf 74.64\pm 0.43$ $\bf 91.68\pm 0.36$
Table 4: Additional experiments on transfer learning. Downstream top-1 classification accuracy of ResNet-50 pretrained via MoCo v2 $+$ SASSL on ImageNet. Accuracy evaluated on six out of twelve target datasets. SASSL generates specialized representations that improve transfer learning performance, both in linear probing and fine-tuning. Mean and standard deviation reported over five random trials.
Target Dataset
ImageNet ImageNet (1%) iNat21 Retinopathy DTD Food101
Linear Probing
None (Default) $72.55\pm 0.67$ $53.23\pm 0.45$ $41.33\pm 0.2$ $\bf 75.88\pm 0.12$ $72.68\pm 0.7$ $73.82\pm 0.1$
ImageNet (Ours) $74.07\pm 0.46$ $56.87\pm 0.43$ $45.01\pm 0.04$ $75.75\pm 0.1$ $73.69\pm 1.22$ $74.43\pm 0.38$
iNat21 (Ours) $74.28\pm 0.38$ $56.76\pm 0.23$ $44.70\pm 0.37$ $75.75\pm 0.17$ $72.75\pm 1.01$ $74.3\pm 0.48$
Retinopathy (Ours) $74.02\pm 0.61$ $\bf 56.99\pm 0.26$ $44.9\pm 0.16$ $75.78\pm 0.08$ $73.73\pm 0.57$ $74.53\pm 0.29$
DTD (Ours) $74.32\pm 0.37$ $56.77\pm 0.36$ $\bf 45.08\pm 0.31$ $75.76\pm 0.11$ $\bf 74.41\pm 1.39$ $\bm{74.88\pm 0.32}$
PBN (Ours) $\bf 74.64\pm 0.43$ $56.9\pm 0.18$ $45.02\pm 0.14$ $75.79\pm 0.07$ $72.77\pm 0.77$ $74.37\pm 0.19$
Fine-tuning
None (Default) $74.89\pm 0.67$ $51.61\pm 0.13$ $77.92\pm 0.14$ $78.89\pm 0.2$ $71.54\pm 0.43$ $87.25\pm 0.11$
ImageNet (Ours) $75.52\pm 0.23$ $51.74\pm 0.14$ $79.21\pm 0.07$ $79.64\pm 0.16$ $\bf 72.31\pm 1.85$ $87.48\pm 0.21$
iNat21 (Ours) $\bf 75.58\pm 0.47$ $\bf 51.86\pm 0.3$ $79.19\pm 0.12$ $79.6\pm 0.23$ $71.35\pm 1.58$ $87.4\pm 0.38$
Retinopathy (Ours) $75.52\pm 0.64$ $51.76\pm 0.26$ $79.23\pm 0.05$ $79.63\pm 0.13$ $72.07\pm 1.61$ $87.39\pm 0.19$
DTD (Ours) $75.24\pm 0.65$ $51.73\pm 0.19$ $\bf 79.24\pm 0.08$ $\bf 79.7\pm 0.15$ $70.59\pm 1.42$ $\bm{87.66\pm 0.23}$
Style Dataset PBN (Ours) $75.05\pm 0.69$ $51.85\pm 0.16$ $79.2\pm 0.1$ $79.63\pm 0.13$ $71.35\pm 0.96$ $87.56\pm 0.38$
@@ A.4 Downstream Training and Testing Settings
A.5 Additional MoCo v2 Results
We complement the MoCo v2 results reported in Tables 4 and 1 by computing both their mean and standard deviation over five random trials.
Downstream performance on ImageNet. Table 3 reports the downstream performance of our SASSL + MoCo v2 representation model, pretrained and linearly probed on ImageNet. Top-1 and top-5 accuracy is computed over five random trials and reported in terms of their mean and standard deviation.
Results show SASSL pretraining and subsequent linear probing on ImageNet yield a notable boost in top-1 classification accuracy, exceeding the default model’s performance by over two standard deviations. This statistically significant improvement underscores the efficacy of incorporating SASSL augmentation into the self-supervised learning process.
Transfer learning performance. Tables 4 and 5 show the transfer learning performance of our SASSL + MoCo v2 model. Linear probing and fine-tuning top-1 accuracy is computed over five random trials and reported in terms of its mean and standard deviation.
Both linear probing and fine-tuning benefit from SASSL pretraining. Notably, linear probing achieves significant gains of up to $10\%$, surpassing default performance by over one standard deviation in most cases, although some target datasets show high variations. While fine-tuning exhibits a smaller improvement gap compared to default models, SASSL representation models consistently outperform baselines by up to $6\%$ in average top-1 accuracy, showcasing their robustness across diverse datasets.
Additionally, the performance differences between different choices of style dataset are generally comparable to the measurement uncertainty, which is typically on the order of 0.3 percentage points. This suggests that the choice of style dataset does not appear to have as significant of an impact as may be expected. It seems like the main benefit is drawn from using a different style rather than due to anything about the style itself. We also show in Section A.6 that using novel styles drawn from natural images does provide a statistically significant improvement compared to synthetic styles.
It is important to highlight that this trend of shrinking improvement gaps between fine-tuning and linear probing occurs with other SSL methods as well. This is because fine-tuning adjusts the entire model to the target dataset, making pretrained weights act mainly as a refined model initialization.
Table 5: Additional experiments on transfer learning. Downstream top-1 classification accuracy of ResNet-50 pretrained via MoCo v2 $+$ SASSL on ImageNet. Accuracy evaluated on six out of twelve target datasets. SASSL generates specialized representations that improve transfer learning performance, both in linear probing and fine-tuning. Mean and standard deviation reported over five random trials.
Target Dataset
CIFAR10 CIFAR100 SUN397 Cars Caltech-101 Flowers
Linear Probing
None (Default) $89.94\pm 0.24$ $71.93\pm 0.48$ $69.96\pm 0.33$ $53.15\pm 0.48$ $88.19\pm 0.75$ $93.39\pm 0.58$
ImageNet (Ours) $90.93\pm 0.28$ $73.26\pm 0.32$ $69.67\pm 0.26$ $\bm{64.87\pm 1.03}$ $89.3\pm 0.23$ $95.27\pm 0.4$
iNat21 (Ours) $\bm{91.04\pm 0.16}$ $73.29\pm 0.28$ $\bm{70.07\pm 0.37}$ $63.96\pm 1.19$ $\bm{89.89\pm 1.07}$ $94.7\pm 0.87$
Retinopathy (Ours) $90.8\pm 0.22$ $73.3\pm 0.38$ $69.63\pm 0.55$ $64.06\pm 0.79$ $89.17\pm 0.28$ $94.94\pm 0.85$
DTD (Ours) $\bf{91.04\pm 0.2}$ $\bm{73.41\pm 0.23}$ $69.71\pm 0.44$ $64.58\pm 0.71$ $89.3\pm 0.26$ $95.24\pm 0.22$
PBN (Ours) $90.85\pm 0.17$ $73.38\pm 0.22$ $69.69\pm 0.26$ $64.12\pm 0.95$ $89.59\pm 0.68$ $\bm{95.45\pm 0.34}$
Fine-tuning
None (Default) $96.91\pm 0.12$ $83.4\pm 0.35$ $74.25\pm 0.13$ $83.63\pm 7.39$ $89.27\pm 0.15$ $95.75\pm 0.24$
ImageNet (Ours) $97\pm 0.09$ $83.21\pm 0.18$ $73.89\pm 0.39$ $\bm{90.33\pm 0.36}$ $88.26\pm 0.35$ $\bm{96.6\pm 0.14}$
iNat21 (Ours) $\bm{97.05\pm 0.14}$ $83.29\pm 0.27$ $74.05\pm 0.3$ $90.04\pm 0.39$ $88.55\pm 0.48$ $95.76\pm 1.75$
Retinopathy (Ours) $96.97\pm 0.09$ $\bm{83.68\pm 0.25}$ $\bm{74.26\pm 0.12}$ $89.96\pm 0.52$ $88.44\pm 0.44$ $96.34\pm 0.28$
DTD (Ours) $96.77\pm 0.41$ $83.28\pm 0.79$ $74.17\pm 0.42$ $89.59\pm 0.59$ $\bm{89.54\pm 1.94}$ $95.59\pm 1.61$
Style Dataset PBN (Ours) $96.97\pm 0.11$ $83.36\pm 0.21$ $74.18\pm 0.46$ $89.75\pm 0.74$ $88.97\pm 0.78$ $95.77\pm 1.78$
A.6 Additional Ablation Studies
SASSL leverages NST to augment pre-training datasets within SSL methods. The proposed technique maintains semantic content by applying transformations solely to the image’s texture. Given a pre-training sample, SASSL treats it as the content image and employs established NST techniques to match the distribution of its low-level features to a chosen style reference. This process creates a new image where the scene’s objects, represented by high-level features (Johnson et al., 2016; Zhang et al., 2018), are preserved while the image’s texture, represented by the distribution of low-level features (Portilla & Simoncelli, 2000; Zhu et al., 2000; Heeger & Bergen, 1995), aligns with the provided style image.
The following section presents a comprehensive exploration of how the properties of the style reference affect SASSL’s generalization to downstream applications.
Effect of the Style Representation in Downstream Performance. We conduct additional experiments to better understand the effect of the style representation $\hat{{\bm{z}}}$ in the downstream task performance of models pretrained via SASSL. Specifically, we replace the style latent code, originally taken from a style image, by (i) i.i.d. Gaussian noise $\hat{{\bm{z}}}\sim{\mathcal{N}}(\bm{\mu},\bm{\Sigma})$, and (ii) the style representation of the content image $\hat{{\bm{z}}}={\bm{z}}_{c}$. Note that latter case is equivalent to using the stylization model ${\mathcal{T}}$ as an autoencoder, since no external style is imposed over the feature maps of the content image.
In both cases, we pretrain and linearly probe a ResNet-50 backbone on ImageNet using MoCo v2 equipped with SASSL. All training and downstream task settings follow our default configurations, as covered in the Appendix Section A.3 and A.4. We also use the recommended blending and interpolation factors $\alpha,\beta\sim{\mathcal{U}}(0.1,0.3)$.
Table 6 shows the linear probing performance obtained by the two scenarios of interest. As reference, we also include the performance of MoCo v2 with the default data augmentation, as well as MoCo v2 via SASSL using an external style dataset (Painter by Numbers). Results show that using noise as style representation boosts top-1 accuracy by $0.24\%$ with respect to the default data augmentation, while using the content as style reference improves performance by $0.8\%$. This implies that using noise as style representation hinders performance with respect to just encoding and decoding the input image via the stylization network $\mathcal{T}$. On the other hand, using an external style dataset boosts up performance by $2.4\%$, which is a significantly larger improvement over the two scenarios of interest.
Results suggest that the style reference has a strong effect on the downstream performance of the pretrained models. Either by replacing the latent representation of a style image by noise or removing the style alignment process and keeping the compression induced by the Stylization network ${\mathcal{T}}$ (by forcing $\hat{{\bm{z}}}={\bm{z}}_{c}$), the improvement provided via Style Transfer data augmentation is significantly smaller than that obtained with our full technique using external style images.
The combined insights from our ablation study and those in Section 5.5 demonstrate that simply incorporating NST into standard augmentations cannot fully account for the observed accuracy gains. Our findings suggest the existence of additional mechanisms that contribute to the delicate balance between the standard augmentation pipeline and NST. These include feature blending, pixel interpolation, feature maps to align, and style references.
Table 6: Effect of the style representation in downstream performance. Different style representation settings are analyzed when pretraining a ResNet-50 backbone via MoCo v2. Performance reported on a single random trial.